The Harvard Family of Bibliography Styles

Peter Williams

May 4, 1993

1 Introduction

This document describes the harvard family of bibliographic styles which are provided in addition to those described in ?) and ?). This style is primarily intended for use with the BibTEX bibliographic database management system. However, provision is also made for hand coding of bibliographies.

2 Citations

There are two primary forms of citation in the harvard style dependent upon whether the reference is used as a noun or parenthetically. Additionally, where there are more than two authors, all authors are listed in the first citation and in subsequent citations just the first author’s name followed by ‘et al.’ is used. The following example from ?) illustrates these points.

The major improvement concerns the structure of the interview (Ulrich & Trumbo 1965, p. 112) . . .. Later reports (Carlson, Thayer, Mayfield & Peterson 1971) record greatly increased interviewer reliability for structured interviews. Wright (1969, p. 408) comments that ‘undoubtedly interviewer skill is directly related to the validity, quantity and quality of the interview output’, and this would suggest some sort of interviewer training is called for. Rowe (1960), for example, found that trained interviewers are better able to evaluate applicants with some measure of reliability. In addition Wexley, Sanders & Yukl (1973) showed that by extensive interviewer training all significant contrast effects could be eliminated. The results of the 1971 study (Carlson et al. 1971) are still relevant, but efforts to . . ..

To facilitate using a citation as a noun a new command \citeasnoun has been created which has the same syntax as the \cite command except that multiple citations are not permitted. The effect of this command is that

As \citeasnoun{btxdoc} and \citeasnoun[Annex~B]{latex} describe \ldots

produces

As ?) and ?, Annex B) describe … whereas

The \BibTeX\ \cite{btxdoc} and \LaTeX\ \cite[Annex~B]{latex} manuals \ldots

produces

The BibTEX (?) and LaTEX (?, Annex B) manuals … Where appropriate, citations are abbreviated automatically after the first

reference when bibliographies are produced by BibTEX. Provision is also made for this feature to be accessed during manual coding.

3 Styles

3.1 Bibliography Styles

There are three bibliography styles currently available within the harvard family, agsm (used in this document) which is based on ?, pp. 95–98), dcu which is based upon the conventions in use in the Design Computing Unit, Department of Architectural and Design Science, University of Sydney and kluwer which aspires to conform to the requirements of Kluwer Academic Publishers. They are invoked by the \bibliographystyle as described in ?, p. 74) and effect the layout of the entries in the bibliography.

3.2 Citation Styles

There are two citation styles currently available within the harvard family, agsm (used in this document) and dcu which for the previous example would produce:

The BibTEX (?) and LaTEX (?, Annex B) manuals … and for multiple citations such as

  The original documentation \cite{btxdoc,latex} say \ldots

the agsm citation style produces

The original documentation (?, ?) say … and the dcu citation style produces

The original documentation (?; ?) say … The default citation style is agsm and both styles have no effect on the appearance of the \citeasnoun citation format.

These styles are invoked by the \citationstyle command, for example:

  \citationstyle{agsm}.

Because these styles affect the format of parenthetical citations, this command should appear before any \cite commands.

4 Doing It By Hand

Hand coding is accomplished much the same as described in ?, p. 73) except that the new command \harvarditem is used in place of \bibitem. The syntax of this command is

\harvarditem [abbr-citation]{full-citation}{citation-year}{cite-key}

where

abbr-citation is the (optional) abbreviated citation (minus the year) to be used in the text subsequent to the first mention of a particular reference,

full-citation is the full citation (minus the year) to be used in the text on the first mention of a particular reference,

citation-year the year portion of the citation including any suffices required to disambiguate citations, and

cite-key is the key used in the \cite and \citeasnoun commands.

5 Acknowledgement

The motivation for this style came from Fay Sudweeks of the Design Computing Unit who also originated the formats for the dcu style and proofread their implementation.